OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles C=30 degrees , c=9, and a=18

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

My book says two but I only got 1 triangle (a right triangle)

OpenStudy (anonymous):

\[\frac{ 9 }{ \sin 30 }=\frac{ 18 }{ sinA }\]

OpenStudy (anonymous):

\[18=\frac{ 18 }{ sinA }\]

OpenStudy (anonymous):

\[sinA=\frac{18}{18}=1\]

OpenStudy (anonymous):

A=90

OpenStudy (primeralph):

That's what I'm getting too.

OpenStudy (anonymous):

30,60,90 triangle, by defintion, only one triangle, but could you argue that with the ambiguous cases of SSA that 180-90=90

OpenStudy (anonymous):

and that additional 90 degrees+30=120 (less than that 180) so you can still make another triangle, but they're both the same triangle.....

OpenStudy (anonymous):

Is my book counting those identical triangles as two triangles??

OpenStudy (primeralph):

Yeah, I know what you mean. sin90 can only give 1 once without blowing the triangle out, so it should be 1 triangle since they ate similar.

OpenStudy (anonymous):

yea....oh well stupid book

OpenStudy (anonymous):

thnx anyways

OpenStudy (primeralph):

Well, let's not say that just yet. Ask your professor though.

OpenStudy (anonymous):

Independence day break .__.

OpenStudy (primeralph):

Oh, cool/

OpenStudy (raden):

one triangle